OpenStudy (anonymous):

In two supplementary angles, the measure of one angle is 6 more than twice the measure of the other. The measures of these two angles are?

OpenStudy (matt101):

Let a and b be the two angles. Since they're supplementary, a + b = 180 --> a = 180 - b According to the second part of the question, a = 2b + 6 Combining the two equations: 180 - b = 2b + 6 3b = 174 b = 58 Since a = 180 - b, a = 180 - 58 = 122 The angles are 58 and 122.